OpenStudy (mayankdevnani):

If centre of circle is \(\large \bf (\pi,e)\),then how many rational points will lie on circle ?

OpenStudy (anonymous):

\[ (x-\pi)^2+(y-e)^2=r^2 \]

OpenStudy (mayankdevnani):

correct,then

OpenStudy (anonymous):

Technically speaking, I suppose \((0,0)\) could lie on the circle with the right radius.

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

\[ x^2-2\pi x+\pi^2+y^2-2e y+e^2 = r^2 \]

OpenStudy (mayankdevnani):

then

OpenStudy (anonymous):

I'm thinking you would analyze how the terms would be in terms of being rational or irrational.

OpenStudy (mayankdevnani):

yeah !!

OpenStudy (mayankdevnani):

and we know that \[\large \bf rational \neq irrational\]

OpenStudy (mayankdevnani):

so answer would be `0`

OpenStudy (anonymous):

Irrational times a rational is irrational.

OpenStudy (mayankdevnani):

correct

OpenStudy (mayankdevnani):

then what do you want to say ? @wio

OpenStudy (dumbcow):

doesn't it depend on whether the radius is rational or not ?

ganeshie8 (ganeshie8):

you have already got one rational point on the circle if you let the circle pass through a rational point to start with

OpenStudy (mayankdevnani):

i don't think radius is irrational

OpenStudy (mayankdevnani):

how ? ganehie

OpenStudy (anonymous):

\[r^2 = (x-\pi)^2+(y-e)^2,~~~~r>0\] I think this makes more sense

ganeshie8 (ganeshie8):

let the circle pass through (0, 0)

OpenStudy (anonymous):

\[ x^2+y^2= r^2 +2\pi x-\pi^2+2e y-e^2 \]It's possible that irrational numbers cancel themselves out.

OpenStudy (mayankdevnani):

ok.. @ganeshie8

OpenStudy (mayankdevnani):

no. @wio

OpenStudy (anonymous):

We could let the radius equal: \[ \sqrt{\pi^2+e^2} \]In this case: \[ (x-\pi)^2+(y-e)^2=\pi^2+e^2 \]And the point \((0,0)\) is on the circle.

ganeshie8 (ganeshie8):

lets see if we can find one more rational point

OpenStudy (mayankdevnani):

oh !! i see

OpenStudy (mayankdevnani):

so (0,0) is one rational point which lies on circle

OpenStudy (mayankdevnani):

correct ??

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

if we pass the circle from point (1,1) then

OpenStudy (anonymous):

I think you can calculate the distance between the centre to the point.

ganeshie8 (ganeshie8):

you don't know yet whether you can pass the same circle through both (0,0) and (1,1) or not

OpenStudy (anonymous):

Giving you the radius

OpenStudy (dumbcow):

hmm any rational point is possible given an irrational radius

OpenStudy (anonymous):

\[ r = \sqrt{(x_1-\pi)^2+(y_1-e)^2} \]You can get any point \(x_1,y_1\) in the circle with the right radius.

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):



OpenStudy (mayankdevnani):

if the point \(\large \bf (x_1,y_1)=(1,1)\),then

OpenStudy (mayankdevnani):

@wio

ganeshie8 (ganeshie8):

for definiteness, lets construct a circle that passes through (0, 0)

ganeshie8 (ganeshie8):

and see if we can find any other rational points on this circle

OpenStudy (mayankdevnani):

no,if we pass the circle through (0,0)

OpenStudy (mayankdevnani):

right ???

OpenStudy (anonymous):

\[ r^2=(1-\pi)^2+(1-e)^2 = (x-\pi)^2+(y-e)^2 \]So \((1,1)\) can be on the circle.

OpenStudy (mayankdevnani):

it will lie on circle or not ??

ganeshie8 (ganeshie8):

(0,0) and (1,1) canot lie on the circle at the same time

OpenStudy (mayankdevnani):

correct !

ganeshie8 (ganeshie8):

but that observation is useless, it wont help us going forward

OpenStudy (mayankdevnani):

i know what i do to move forward as @wio tells me ! @ganeshie8

OpenStudy (mayankdevnani):

*told

ganeshie8 (ganeshie8):

idk what to do yet, im still thinking. what do you have in mind ?

OpenStudy (mayankdevnani):

now my question is this :- \[\large \bf if~we~pass~circle~through~(1,1)~instead~of(0,0),~then\]

ganeshie8 (ganeshie8):

then nothign happens, its the same story :O

OpenStudy (anonymous):

You can use the distance formula and check for yourself

OpenStudy (mayankdevnani):

@ganeshie8 if circle pass through (0,0),then \[\large \bf (x- \pi)^2+(y-e)^2=r^2\]

ganeshie8 (ganeshie8):

you got one rational point on ur circle and you're stuck at the same place

OpenStudy (mayankdevnani):

plug (0,0), \[\large \bf r^2=\pi^2+e^2\]

OpenStudy (mayankdevnani):

\[\large \bf \pi^2+e^2=r^2\] \[\large \bf \pi^2+e^2=\pi^2+e^2\] so we can say that (0,0) is rational point which lie on circle

OpenStudy (mayankdevnani):

@ganeshie8 and @wio